Steven purchases a car for$90,000. If the value of the car decreases per year, what is the average rate of decrease if the value of the car is $30,000 after 8 years?
Answer by stanbon(75887)

Steven purchases a car for$90,000. If the value of the car decreases per year, what is the average rate of decrease if the value of the car is $30,000 after 8 years?
-----
Solve:
30,000 = x^8*90000
----
x^8 = 1/3
----
x = (1/3)^(1/8)
-----
x = 0.8717
-----
Average decrease = 1-0.8717 = 0.1283 = 12.83%
------------------------
Cheers,
Stan H.